### v2

# Latent Functional Maps

Abstract
Neural models learn data representations that lie on low-dimensional manifolds, yet modeling the relation between these representational spaces is an ongoing challenge. By integrating spectral geometry principles into neural modeling, we show that this problem can be better addressed in the functional domain, mitigating complexity, while enhancing interpretability and performances on downstream tasks. To this end, we introduce a multi-purpose framework to the representation learning community, which allows to: (i) compare different spaces in an interpretable way and measure their intrinsic similarity; (ii) find correspondences between them, both in unsupervised and weakly supervised settings, and (iii) to effectively transfer representations between distinct spaces. We validate our framework on various applications, ranging from stitching to retrieval tasks, demonstrating that latent functional maps can serve as a swiss-army knife for representation alignment.
**footnotetext: Equal Contributionfootnotetext: Work done while visiting ISTA
1 Introduction
Neural Networks (NNs) learn to represent high-dimensional data as elements of lower-dimensional latent spaces. While much attention is given to the model’s output for tasks such as classification, generation, reconstruction, or denoising, understanding the internal representations and their geometry is equally important. This understanding facilitates reusing representations for downstream tasks [48] and the comparison between different models [14, 29], thereby broadening the applicability of NNs, and understanding their structure and properties. Moreover, recent studies have shown that distinct neural models often develop similar representations when exposed to similar stimuli, a phenomenon observed in both biological [9, 16] and artificial settings [28, 14, 29]. Notably, even with different initializations of the same neural architecture, internal representations of distinct models can often be aligned through a linear transformation [47, 38]. This indicates a level of consistency in how NNs process information, highlighting the importance of exploring and understanding these internal representations.
In this paper, we shift our focus from characterizing relationships between samples in distinct latent spaces to modeling a map among function spaces defined on these latent manifolds. To this end, we propose leveraging spectral geometry principles by applying the framework of functional maps [34] to the field of representation learning. Functional maps represent correspondences between function spaces on different manifolds, providing a new perspective on the problem of latent space communication. In this setting, many difficult constraints can be easily manipulated and expressed compactly [35]. For instance, as shown in Figure 1, the mapping in the functional space ($C$) becomes a linear map with a sparse structure.
Originally used in 3D geometry processing and more recently on graphs applications [46, 36], we extend this framework to handle arbitrary large dimensional latent spaces. Our proposed method, Latent Functional Map (LFM), is a flexible tool that allows (i) to compare distinct representational spaces, (ii) to find correspondences between them both in an unsupervised and weakly supervised way, and (iii) to transfer information between them.
[…]
Figure 1: Framework overview: given two spaces $\mathcal{X}$, $\mathcal{Y}$ their samples lie on two manifolds $\mathcal{M}$, $\mathcal{N}$, which can be approximated with the KNN graphs $G_{\mathcal{X}}$,$G_{\mathcal{Y}}$; We can optimize for a latent functional map $C$ between the eigenbases of operators defined on the graphs. This map serves as a map between functions defined on the two manifolds and can be leveraged for (i) comparing representational spaces, (ii) solving correspondence problems, and (iii) transferring information between the spaces.
Our contributions can be listed as follows:
  • We introduce the framework of Latent Functional Maps as a way to model the relation between distinct representational spaces of neural models.
  • We show that LFM allows us to find correspondences between representational spaces, both in weakly supervised and unsupervised settings, and to transfer representations across distinct models.
  • We showcase LFM capabilities as a meaningful and interpretable similarity measure between representational spaces.
  • We validate our findings in retrieval and stitching tasks across different models, modalities and datasets, demonstrating that LFMs can lead to better performance and sample efficiency than other methods.
2 Related Work
Similarity between latent spaces
Comparing representations learned by neural models is of fundamental importance for a diversity of tasks, ranging from representation analysis to latent space alignment and neural dynamics. In order to do so, a similarity measure between different spaces must be defined [13]. This can range from functional similarity (matching the performance of two models) to similarity defined in representational space [14], which is where our framework falls in. A classical statistical method is Canonical Correlation Analysis (CCA) [11], known for its invariance to linear transformations. Various adaptations of CCA aim to enhance robustness, such as Singular Vector Canonical Correlation Analysis (SVCCA) [37], or to decrease sensitivity to perturbations using methods like Projection-Weighted Canonical Correlation Analysis (PWCCA) [28]. Closely related to these approaches, Centered Kernel Alignment (CKA) [14] measures the similarity between latent spaces while ignoring orthogonal transformations. However, recent research [7] reveals that CKA is sensitive to local shifts in the latent space.
We propose to leverage LFMs as a tool to measure the similarity, or how much two spaces differ from an isometry w.r.t. to the metric that has been used to construct the graph.
Latent communication
This relatively new concept, introduced by [29], builds on the hypothesis that latent spaces across neural networks (pre-)trained with many variation factors, from random seed initialization to architecture or even data modality, are intrinsically compatible. This notion is supported by numerous empirical studies [28, 20, 14, 4, 42, 1, 45, 18, 19, 2, 30, 6], with the phenomenon being particularly evident in large and wide models [39, 23]. The core idea is that relations between data points (i.e., distances according to some metric) are preserved across different spaces because the high-level semantics of the data are the same and neural networks learn to encode them similarly [12]. With this "relative representation", the authors show that it is possible to stitch [19] together model components coming from different models, with little to no additional training, as long as a partial correspondence of the spaces involved is known. Indeed, [17, 22, 25, 27] show that a simple linear transformation is usually enough to effectively map one latent space into another, measuring the mapping performance via desired downstream tasks.
With LFMs, we change the perspective from merely relating samples of distinct latent spaces to relating function spaces defined on the manifold that the samples approximate, showing that processing information in this dual space is convenient as it boosts performance while also being interpretable.
Functional Maps.
The representation we propose is directly derived from the functional maps framework for smooth manifolds introduced in the seminal work by [34]. This pioneering study proposed a compact and easily manipulable mapping between 3D shapes. Subsequent research has aimed at enhancing this framework. For instance, [31] introduced regularization techniques to improve the informativeness of the maps, while [24] developed refinement methods to achieve more accurate mappings. The functional map framework has been extended as well outside the 3d domain, for example, in [46] and [10], who applied the functional framework to model correspondences between graphs, and in [36], who demonstrated its utility in graph learning tasks. In particular, they have shown that the functional map representation retains its advantageous properties even when the Laplace basis is computed on a graph.
Inspired by these advancements, our work leverages the functional representation of latent spaces of neural models. We demonstrate how this representation can be easily manipulated to highlight similarities and facilitate the transfer of information between different spaces, thereby extending the applicability of the functional maps framework to the domain of neural latent spaces.
3 Background
This section provides the basic notions to understand the framework of functional maps applied to manifolds. We refer to [35] for a comprehensive overview.
Consider two manifolds $\mathcal{M}$ and $\mathcal{N}$ equipped with a basis such that any function $f:\mathcal{M}\rightarrow\mathbb{R}$ can be represented as a linear combination of basis functions $\mathbf{\Phi}_{\mathcal{M}}$: $f=\sum_{i}a_{i}\Phi_{i}^{\mathcal{M}}=\mathbf{a}\mathbf{\Phi}_{\mathcal{M}}$. Given the correspondence $T:\mathcal{M}\rightarrow\mathcal{N}$ between points on these manifolds, for any real-valued function $f:\mathcal{M}\rightarrow\mathbb{R}$, one can construct a corresponding function $g:\mathcal{N}\rightarrow\mathbb{R}$ such that $g=f\circ T^{-1}$. In other words, the correspondence $T$ defines a mapping between two function spaces $T_{F}:\mathcal{F}(\mathcal{M},\mathbb{R})\rightarrow\mathcal{F}(\mathcal{N},\mathbb{R})$. [34] showed how such a mapping is linear and can be represented as a (possibly infinite) matrix $\mathbf{C}$ such that for any function $f$ represented as a vector of coefficients $\mathbf{a}$, we have $T_{F}(\mathbf{a})=\mathbf{C}\mathbf{a}$.
The functional representation is particularly well-suited for map inference (i.e., constrained optimization). When the underlying map $T$ (and by extension the matrix $\mathbf{C}$) is unknown, many natural constraints on the map become linear constraints in its functional representation. In practice, the simplest method for recovering an unknown functional map is to solve the following optimization problem:
$\underset{\mathbf{C}}{\arg\min}||\mathbf{C}\mathbf{A}-\mathbf{B}||^{2}+\rho(\mathbf{C})$ (1)
where $\mathbf{A}$ and $\mathbf{B}$ are sets of corresponding functions expressed in the bases on $\mathcal{M}$ and $\mathcal{N}$, respectively, and $\rho(\mathbf{C})$ represents additional constraints deriving from the properties of the matrix $\mathbf{C}$ [35]. When the shapes are approximately isometric and the descriptors are well-preserved by the (unknown) map, this procedure provides a good approximation of the underlying map. In the case where the correspondence $T$ is encoded in a matrix $\mathbf{S}$, the functional map can be retrieved as $\mathbf{C}=\mathbf{\Phi}_{\mathcal{N}}^{{\dagger}}\mathbf{S}\mathbf{\Phi}_{\mathcal{M}}$ where $\mathbf{\Phi}_{\mathcal{M}}$ and $\mathbf{\Phi}_{\mathcal{N}}$ are the bases of the functional spaces $\mathcal{F}(\mathcal{M},\mathbb{R})$ and $\mathcal{F}(\mathcal{N},\mathbb{R})$, respectively, and denotes the pseudo-inverse.
4 Method
4.1 Setting
We consider deep neural networks $f:=f_{1}\circ f_{2}\circ...f_{n}$ where each layer $f_{i}$ is associated to a representational space $\mathcal{X}$ corresponding to the image of $f_{i}$. We assume that elements $x\in\mathcal{X}$ are sampled from a latent manifold $\mathcal{M}$. Considering pairs of spaces $(\mathcal{X},\mathcal{Y})$, and corresponding manifolds $\mathcal{M},\mathcal{N}$ our objective is to characterize the relation between them by mapping the space of functions $\mathcal{F}(\mathcal{M})$ to $\mathcal{F}(\mathcal{N})$. Our framework is depicted in Figure 1. In the following, we will start by approximating $\mathcal{X}$ from a sample estimate, building a graph in the latent space.
Building the graph
To leverage the geometry of the underlying manifold, we model the latent space of a neural network building a symmetric k-nearest neighbor (k-NN) graph [21]. Given a set of samples $X=\{x_{1},\ldots,x_{n}\}$, we construct an undirected weighted graph $G=(X,E,\mathbf{W})$ with nodes $X$, edges $E$, and weight matrix $\mathbf{W}$. The weight matrix is totally characterized by the choice of distance function $d(x_{i},x_{j})$ with $x_{i},x_{j}\in X$. Suitable choices include the L2 metric or the angular distance. Edges $E$ are defined as $E=\{(x_{i},x_{j})\in X\times X\mid x_{i}\sim_{\text{k}}x_{j}\text{ or }x_{j}\sim_{\text{k}}x_{i}\}$, where $x_{i}\sim_{\text{k}}x_{j}$ indicates that $x_{j}$ is among the $k$ nearest neighbors of $x_{i}$. The weight matrix $\mathbf{W}\in\mathbb{R}^{n\times n}_{\geq 0}$ assigns a weight $\omega(x_{i},x_{j})$ to each edge $(x_{i},x_{j})\in E$, and $\mathbf{W}_{i,j}=0$ otherwise.
Next, we define the associated weighted graph Laplacian $\mathcal{L}_{G}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{W}\mathbf{D}^{-1/2}$, where $\mathbf{D}$ is the diagonal degree matrix with entries $\mathbf{D}_{i,i}=\sum_{j=1}^{n}\mathbf{W}_{i,j}$. $\mathcal{L}_{G}$ is a positive semi-definite, self-adjoint operator [44]), therefore, it admits an eigendecomposition $\mathcal{L}_{G}=\mathbf{\Phi}_{G}\mathbf{\Lambda}_{G}\mathbf{\Phi}_{G}^{T}$, where $\mathbf{\Lambda}_{G}$ is a diagonal matrix containing the eigenvalues, and $\mathbf{\Phi}_{G}$ is a matrix whose columns are the corresponding eigenvectors. The eigenvectors form an orthonormal basis for the space of functions defined on the graph nodes (i.e., $\mathbf{\Phi}_{G}^{T}\mathbf{\Phi}_{G}=\mathbf{I}$).
Throughout this paper, we assume the eigenvalues (and corresponding eigenvectors) are sorted in non-descending order $0=\Lambda_{1}\leq\Lambda_{2}\leq\dots\leq\Lambda_{n}$. One may consider a subset of eigenvectors, namely those associated with the $k$ smallest eigenvalues, to compactly approximate a graph signal, employing techniques akin to Fourier analysis.
As demonstrated in many recent works [41, 5], the eigenvalues and eigenvectors of the graph Laplacian associated with a k-NN graph approximate the weighted Laplace-Beltrami operator, placing us in a setting similar to the original one of [34].
4.2 Latent Functional Maps
We consider pairs of latent spaces $\mathcal{X}$ and $\mathcal{Y}$. We model each space using a subset of training samples $X=\{x_{1},\ldots,x_{n}\}$ and $Y=\{y_{1},\ldots,y_{n}\}$ and build the k-NN graphs $G_{X}$ and $G_{Y}$ from these samples, respectively. For each graph, we compute the graph Laplacian $\mathcal{L}_{G}$ and derive the first $k$ eigenvalues $\Lambda_{G}$ and eigenvectors $\mathbf{\Phi}_{G}=[\phi_{1},\ldots,\phi_{k}]$, which serve as the basis for the function space defined on the latent spaces.
Given the set of corresponding functions $\mathbf{F}_{G_{X}}=[f^{G_{X}}_{1},\ldots,f^{G_{X}}_{n_{f}}]$ and $\mathbf{F}_{G_{Y}}=[f^{G_{Y}}_{1},\ldots,f^{G_{Y}}_{n_{f}}]$, we consider the optimization problem defined in Equation 1 and incorporate regularizers for Laplacian and descriptor operator commutativity, as defined in [31]:
$\underset{\mathbf{C}}{\arg\min}||\mathbf{C}\hat{\mathbf{F}}_{G_{X}}-\hat{\mathbf{F}}_{G_{Y}}||^{2}+\alpha\rho_{\mathcal{L}}(\mathbf{C})+\beta\rho_{f}(\mathbf{C})$ (2)
where $\hat{\mathbf{F}}_{G}=\mathbf{\Phi}_{G}^{T}\mathbf{F}_{G}$ are the spectral coefficients of the functions $\mathbf{F}_{G}$, $\rho_{\mathcal{L}}$ and $\rho_{f}$ are the Laplacian and descriptor operator commutativity regularizers respectively. We specify how we compute the regularizers in Appendix A.
As a set of corresponding functions, we use the geodesic distance functions computed from a point $x\in X$ to all other points in $X$, where $x$ is a point for which we know the corresponding point $y\in Y$ in the other latent space. Once we have solved the optimization problem defined in Equation 2, we refine the resulting functional map $\mathbf{C}$ using the algorithm proposed by [24].
4.3 LFMs as a similarity measure
Once computed, the functional map $\mathbf{C}$ can serve as a measure of similarity between spaces. The reason is that for isometric transformations between manifolds, the functional map is volume preserving (see Thm 5.1 in [34]), and this is manifested in orthogonal $\mathbf{C}$. By defining the inner product between functions $h_{1},h_{2}\in\mathcal{F}({M})$ as $\langle h_{1},h_{2}\rangle=\int_{\mathcal{M}}h_{1}(x)h_{2}(x)\mu(x)$, it holds that $\langle h_{1},h_{2}\rangle=\langle\hat{h}_{1},\hat{h}_{2}\rangle$ when the map preserves the local area, where $\hat{h}$ denotes the functional representation of $h$. In other words, when the transformation between the two manifolds is an isometry, the matrix $\mathbf{C}^{T}\mathbf{C}$ will be diagonal. By measuring the ratio between the norm of the off-diagonal elements of $\mathbf{C}^{T}\mathbf{C}$ and the norm of its diagonal elements, we can define a measure of similarity $sim(X,Y)=1-\frac{||\text{off}((\mathbf{C}^{T}\mathbf{C})||_{F}}{||\text{diag}(\mathbf{C}^{T}\mathbf{C})||_{F}}$. Furthermore, this quantity is interpretable; the first eigenvector of $\mathbf{C}^{T}\mathbf{C}$ can act as a signal to localize the area of the target manifold where the map has higher distortion [33].
4.4 Transfering information with LFM
The functional map computed between two latent spaces can be utilized in various ways to transfer information from one space to the other. In this paper, we focus on two methods: (i) Expressing arbitrary points in the latent space as distance function on the graph and transferring them through the functional domain; (ii) Obtaining a point-to-point correspondence between the representational spaces from the LFM, starting from none to few known pairs, and leverage off-the-shelf methods to learn a transformation between the spaces. Additional strategies could be explored in future work.
Space of Functional Coefficients
The space of functional coefficients offers an alternative representation for points in the latent space $\mathcal{X}$. Using the equation $\hat{f}_{G}=\mathbf{\Phi}_{G}^{T}f_{G}$, any function $f_{G}\in\mathcal{F}(G,\mathbb{R})$ can be uniquely represented by its functional coefficients $\hat{f}_{G}$. We leverage this property to represent any point $x\in\mathcal{X}$ as a distance function $f_{d}\in\mathcal{F}(G,\mathbb{R})$ from the set of points $X_{G}$, which correspond to the nodes of the graph $G$. The functional map $\mathbf{C}$ between two latent spaces $\mathcal{X}$ and $\mathcal{Y}$ aligns their functional representations, enabling the transfer of any function from the first space to the second. This functional alignment can be used similarly to the method proposed by [29] to establish a "relative" space where the representational spaces $\mathcal{X}$ and $\mathcal{Y}$ are aligned.
Extending spaces correspondences
As explained in Section 3, the functional map $\mathbf{C}$ represents the bijection $T$ in a functional form. [34] demonstrated that this bijection can be retrieved as a point-to-point map by finding the nearest neighbor for each row of $\mathbf{\Phi}_{G_{Y}}\mathbf{C}$ in $\mathbf{\Phi}_{G_{X}}$. This process can be efficiently implemented using algorithms such as kd-tree. Given a few correspondences (anchors) between the two spaces $\mathcal{X}$ and $\mathcal{Y}$, we can extend these correspondences to the entire set of nodes $X$ and $Y$. This extended set of anchors can then be used to determine a transformation between the latent spaces, as described by [22]. In the following section, we demonstrate that by using a small number of anchors (typically $\leq 50$), we can retrieve optimal transformations that facilitate near-perfect stitching and retrieval.
5 Experiments
5.1 Analysis
We demonstrate the benefits of using latent functional maps for comparing distinct representational spaces, using the similarity metric defined in Section4.3
[…]
Figure 2: Similarity across layers Similarity matrices between internal layer representations of CIFAR10 comparing our LFM-based similarity with the CCA and CKA baselines, averaged across 10 models. For each method, we report the accuracy scores for matching the corresponding layer by maximal similarity.
Experimental setting
In order to validate experimentally if LFMs can serve as a good measure of similarity between distinct representational spaces, we run the same sanity check as in [14]. We train 10 CNN models (the architecture is depicted in Appendix B.1) on the CIFAR-10 dataset [15], changing the initialization seed. We compare their inner representations at each layer, excluding the logits and plot them as a similarity matrix, comparing with Central Kernel Alignment (CKA) measure [14] and Canonical Correlation Analysis (CCA) [11, 37]. We then measure the accuracy of identifying corresponding layers across models and report the results comparing with CKA and CCA as baselines. For CCA, we apply average pooling on the spatial dimensions to the embeddings of the internal layers, making it more stable numerically and boosting the results for this baseline compared to what was observed in [14].
Result analysis
Figure 2 shows that our LFM-based similarity measure behaves correctly as CKA does. Furthermore, the similarities are less spread around the diagonal, favoring a slightly higher accuracy score in identifying the corresponding layers across models.
While CKA (Centered Kernel Alignment) is a widely used similarity metric in deep learning, recent research by [7] has shown that it can produce unexpected or counter-intuitive results in certain situations. Specifically, CKA is sensitive to transformations that preserve the linear separability of two spaces, such as local translations. Our proposed similarity measure is robust to these changes and demonstrates greater stability compared to CKA.
Experimental setting
We compute the latent representations from the pooling layer just before the classification head for the CIFAR10 train and test sets. Following the setup in [7], we train a Support Vector Machine (SVM) classifier on the latent representations of the training samples to find the optimal separating hyperplane between samples of one class and others. We then perturb the samples by translating them in a direction orthogonal to the hyperplane, ensuring the space remains linearly separable. We measure the CKA and LFM similarities as functions of the perturbation vector norm, as shown in Figure 3(a). In the accompanying plot on the right, we visualize the area distortion of the map by projecting the first singular component of the LFM $\mathbf{C}$ into the perturbed space and plotting it on a 2d TSNE [43] projection of the space.
Result Analysis
We start by observing that when the latent space is perturbed in a way that still preserves its linear separability, it should be considered identical from a classification perspective, as this does not semantically affect the classification task. Figure 3(a) shows that while CKA degrades as a function of perturbation intensity, the LFM similarity remains stable to high scores. To understand this difference, we can visualize the area distortion as a function of the samples by projecting the first singular component of $\mathbf{C}$ onto the perturbed space. In Figure 3(b), we use t-SNE [43] to project the perturbed samples and the distortion function into 2D. The visualization reveals that distortion is localized to the samples corresponding to the perturbed class.
[…]
(a) CKA vs LFM similarity
[…]
(b) Localized area distortion
Figure 3: Robustness of LFM similarity Left: Similarity scores as a function of perturbation strength: while the CKA baseline degrades, our LFM similarity scores are robust to perturbations that preserve linear separability of the space. Right: Visualization of area distortion of the map by projecting the first singular component of the LFM in the perturbed space: the distortion localizes on the samples of the perturbed class, making LFM similarity interpretable.
5.2 Zero-shot stitching
We test the use of the latent functional map in the task of zero-shot stitching, as defined in [29], to combine independent encoders and decoders (e.g., classifiers, generators) without subsequent training or fine-tuning.
[…]
(a) Coarse labelling
[…]
(b) Fine labelling
Figure 4: Stitching on CIFAR100.Accuracy performance in stitching between image encoders trained on CIFAR100, comparing orthogonal transformation (Ortho) and LFM+Ortho at varying anchor counts. Also shown is LFM+Ortho (Labels), which uses the dataset labels instead of anchors. Results are presented for (a) coarse and (b) fine-grained labeling, with mean accuracy values reported on each box.
Experimental Setting
We consider four pre-trained image encoders (see Appendix B.2 for details) and stitch their latent spaces to perform classification using a Support Vector Machine (SVM) on two different labelings of CIFAR100 [15]: coarse and fine-grained. To evaluate the effectiveness of integrating the functional map, we extend the correspondences to determine an orthogonal transformation [22] between the latent spaces. For each encoder, we compute a graph of 3,000 points with 300 neighbors per node. We optimize the problem in Equation 2 using the first 50 eigenvectors of the graph Laplacian and consider two different descriptors: the distance functions defined from the anchors (LFM+Ortho) and the labels (LFM+Ortho (Labels)). For each dataset class, the latter provides an indicator function with 1 if the point belongs to the class and 0 otherwise. This descriptor type does not require any anchor as input, representing a pioneering example of stitching requiring no additional information beyond the dataset.
Result Analysis
Figure 4 presents the accuracy results for all possible combinations of encoder stitching. The addition of the latent functional map (LFM+Ortho) shows higher performance with a low number of anchors in both labelings of CIFAR100. Even without any anchors, the label descriptors (LFM+Ortho (Labels)) provide the best performance for the latent functional map framework in both labelings. Computing the orthogonal transformation directly from the anchors (Ortho) proves to have comparable performance only with 500 anchors, where the performance of LFM is limited by the number of eigenvectors used.
This experiment shows that the latent functional map is highly effective when few anchors are available ($\leq 50$). It significantly enhances performance in zero-shot stitching tasks, outperforming direct orthogonal transformations at low or no anchor counts. This suggests that the latent functional map method provides a robust means of aligning latent spaces with minimal correspondence data, making it a valuable tool for tasks requiring the integration of independently trained models.
5.3 Retrieval
We extend our analysis to the retrieval task, where we look for the most similar embedding in the aligned latent space.
[…]
(a) MRR score at increasing number of anchors
[…]
(b) Latent space
[…]
(c) Functional space
Figure 5: Retrieval of word embeddings. We compare the retrieval performance of the functional map framework with state-of-the-art models as the number of anchors increases. The left panel shows the Mean Reciprocal Rank (MRR) across different numbers of anchors. The right panels depict the first two components of PCA for a subsample of the latent space (b) and the functional space (c), both before and after alignment using the functional map.
Experimental Setting
We consider two different English word embeddings, FastText [3] and Word2Vec [26]. Following the approach of [29], we extract embeddings of 20K words from their shared vocabulary using pre-trained models. We use 2K random corresponding samples to construct the k-NN graphs and evaluate the retrieval performance on the remaining 18K word embeddings. We test two settings in our experiments:
  1. 1.
    Aligning functional coefficients (LFM Space).
  2. 2.
    Computing an orthogonal transformation using the correspondences obtained by the functional map (LFM+Ortho).
For this experiment, we construct k-NN graphs with a neighborhood size of 300 and compute the functional map using the first 50 eigenvectors. We evaluate the methods’ performance using the Mean Reciprocal Rank (MRR), as detailed in Appendix B.4. Our functional map methods are compared with the method proposed by [29] (Relatives) and the orthogonal transformation method proposed by [22] (Ortho).
Result Analysis
Figure 5 shows the performance of these methods as the number of anchors increases. The numerical results are detailed in Table 2 in the Appendix. The functional map significantly improves performance with just 5 anchors, achieving an MRR of over 0.8. As the number of anchors increases, the performance of competing methods improves but still falls short of FMAP+Transform at 300 anchors, which reaches an MRR of 0.99. Interestingly, the performance of the functional map methods does not improve beyond 5 anchors, suggesting that this number of anchors is sufficient to achieve an optimal functional map between the spaces. In Appendix C.1, we further analyze how the results improve as the number of eigenvectors used to compute the functional map increases. Notably, the MRR score drastically increases to over 0.6 with more than 25 eigenvectors.
These results confirm that the latent functional map is a valuable tool in settings with little knowledge about correspondences. It significantly enhances retrieval performance with a minimal number of anchors, making it an efficient approach for aligning latent spaces. Moreover, its performance can be improved using a higher number of eigenvectors.
6 Conclusions
In this paper, we explored the application of latent functional maps (LFM) to better understand and utilize the internal representations of neural networks. Our approach leverages the principles of spectral geometry to model and analyze the relationships between different latent spaces, providing a robust framework for comparing and aligning representations across various models. By extending the functional map framework to high-dimensional latent spaces, we offer a versatile method for comparing distinct representational spaces, finding correspondences between them in both unsupervised and weakly supervised settings (few or no anchors) and transferring information across different models.
Limitations and future work
The performance of LFM can be sensitive to the number of eigenvectors used, as seen in some of our experiments. Finding the optimal number of eigenvectors without extensive experimentation remains an open problem. The current framework may not handle very complex transformations between latent spaces as effectively as simpler linear transformations. Expanding the framework to manage more complex transformations is an area for future exploration. Future research can further explore the potential of LFM in other domains and tasks, and modalities. The versatility of the functional representation can be further explored to define new ad-hoc constraints and regularizers for different settings. In particular, ts performance in fully unsupervised settings (with no anchors) is still an area requiring further research and improvement.
In conclusion, the latent functional map framework offers a novel and effective approach to understanding and utilizing neural network representations, promising significant advancements in both theoretical and practical aspects of machine learning.
Appendix A Latent Functional Map
A.1 Additional Regularizers
In Equation 2, we improve the computation of the functional map by incorporating two additional regularizers: Laplacian commutativity and descriptor operator commutativity. Both regularizers exploit the preservation of linear functional operators $\mathbf{S}^{G}:\mathcal{F}(G,\mathbb{R})\rightarrow\mathcal{F}(G,\mathbb{R})$, enforcing that the functional map $\mathbf{C}$ commutes with these operators: $||\mathbf{S}^{G}_{i}\mathbf{C}-\mathbf{C}\mathbf{S}^{G_{X}}_{i}||=0$.
The Laplacian commutativity regularizer, first introduced by [34], is formulated as:
$\rho_{\mathcal{L}}(\mathbf{C})=||\mathbf{\Lambda}_{G_{Y}}\mathbf{C}-\mathbf{C}\mathbf{\Lambda}_{G_{X}}||^{2}$ (3)
where $\mathbf{\Lambda}_{G}$ represents the diagonal matrices of eigenvalues. This regularizer ensures that the functional map $\mathbf{C}$ preserves the spectral properties of the Laplacian.
The descriptor operator commutativity regularizer, introduced by [31], extracts more detailed information from a given descriptor, resulting in a more accurate functional map even with fewer descriptors. The formulation of this regularizer is as follows:
$\rho_{f}(\mathbf{C})=\sum_{i}||\mathbf{S}^{G_{Y}}_{i}\mathbf{C}-\mathbf{C}\mathbf{S}^{G_{X}}_{i}||^{2}$ (4)
where $\mathbf{S}^{G}_{i}=\mathbf{\Phi}_{G}^{T}\text{Diag}(f^{G}_{i})\mathbf{\Phi}_{G}$ are the descriptor operators.
Appendix B Experimental details
B.1 Architecture Details
All non-ResNet architectures are based on All-CNN-C [40]
Tiny-10
$3\times 3$ conv. 16-BN-ReLu ×$\times 2$
$3\times 3$ conv. 32 stride 2-BN-ReLu
$3\times 3$ conv. 32-BN-ReLu ×$\times 2$
$3\times 3$ conv. 64 stride 2-BN-ReLu
$3\times 3$ conv. 64 valid padding-BN-ReLu
$1\times 1$ conv. 64-BN-ReLu
Global average pooling
Logits
Table 1:
B.2 Pre-trained models
In Section 5.2 we used four pretrained models: 3 variations of [8] (’google-vit-base-patch16-224’, ’google-vit-large-patch16-224’, ’WinKawaks-vit-small-patch16-224’) and the model proposed by [32] ( ’facebook-dinov2-base’).
B.3 Parameters and resources
In all our experiments we used gpu rtx 3080ti and 3090. In order to compute the eigenvector and functional map on a graph of 3k nodes we employ not more than 2 minutes.
B.4 Mean Reciprocal Rank (MRR)
Mean Reciprocal Rank (MRR) is a commonly used metric to evaluate the performance of retrieval systems [29]. It measures the effectiveness of a system by calculating the rank of the first relevant item in the search results for each query.
To compute MRR, we consider the following steps:
  1. 1.
    For each query, rank the list of retrieved items based on their relevance to the query.
  2. 2.
    Determine the rank position of the first relevant item in the list. If the first relevant item for query $i$ is found at rank position $r_{i}$, then the reciprocal rank for that query is $\frac{1}{r_{i}}$.
  3. 3.
    Calculate the mean of the reciprocal ranks over all queries. If there are $Q$ queries, the MRR is given by:
    $\text{MRR}=\frac{1}{Q}\sum_{i=1}^{Q}\frac{1}{r_{i}}$ (5)
    Here, $r_{i}$ is the rank position of the first relevant item for the $i$-th query. If a query has no relevant items in the retrieved list, its reciprocal rank is considered to be zero.
MRR provides a single metric that reflects the average performance of the retrieval system, with higher MRR values indicating better performance.
Appendix C Additional Results
C.1 Retrieval
Table 2: MRR Score for the retrieval of word embeddings. We report the value of the results depicted in Figure 5 adding more kind transformation between spaces (Orthogonal, Linear and Affine).
Number of anchors
Method 2 5 10 25 50 75 100 150 200 300
LFM+Ortho 0.01 0.99 0.99 0.99 0.99 0.99 0.99 0.99 0.99 0.99
LFM+Linear 0.01 0.99 0.99 0.99 0.99 0.99 0.99 0.99 0.99 0.99
LFM+Affine 0.01 0.99 0.99 0.99 0.99 0.99 0.99 0.99 0.99 0.99
Ortho 0.01 0.01 0.01 0.03 0.15 0.34 0.60 0.82 0.93 0.97
Linear 0.01 0.01 0.01 0.05 0.26 0.49 0.66 0.77 0.74 0.01
Affine 0.01 0.01 0.01 0.04 0.19 0.45 0.64 0.81 0.89 0.95
LFM Space 0.01 0.85 0.85 0.85 0.85 0.85 0.85 0.85 0.85 0.85
Relatives 0.01 0.01 0.05 0.28 0.55 0.72 0.79 0.84 0.87 0.90
[…]
Figure 6: Retrieval of word embeddings for increasing number of eigenvectors. We report the MMR score for the methods in Figure 5 with fixed number of anchors (5) and increasing number of eigenvectors.
In Table 2, we report the numerical results for the experiment in Figure 5 adding more transformations from the method of [22]: orthogonal (Ortho), linear (Linear) and affine (Affine). From the value in the table, we can see that all the methods that involve the latent functional map (LFM) saturate at 5 anchors, reaching top performance. In Figure 6, we show how the performance of the latent functional map methods depends on the number of eigenvectors used to compute the map. In particular, we notice that the performance drastically increases at 25 eigenvectors, reaching the same score when using the functional map computed from the ground truth correspondences (LFMGT).